# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v2

@@ A.6.3 The $\textsc{NN}_{\textsc{sum}}$ model is Lipschitz continuous with respect to the matching distance
$\bigg{|}\bigg{|}f_{\text{MLP}_{2}}\bigg{(}\sum_{i=1}^{M}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg{)}-f_{\text{MLP}_{2}}\bigg{(}\sum_{j=1}^{M}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg{)}\bigg{|}\bigg{|}$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{M}(X,Y)$
@@
A.6.4 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to the matching distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $X\in\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$, $\forall X\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. Let $\pi^{*}$ denote the matching produced by the solution of the matching distance function:
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{M}}\sum_{i=1}^{M}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}||\$
@@ A.6.4 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to the matching distance
$\bigg{|}\bigg{|}f_{\text{MLP}_{2}}\Big{(}\big{(}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\big{)}_{\text{max}}\Big{)}-f_{\text{MLP}_{2}}\Big{(}\big{(}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\big{)}_{\text{max}}\Big{)}\bigg{|}\bigg{|}$ $\leq\text{Lip}(f_{\text{MLP}_{2}})\Big{|}\Big{|}\big{(}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\big{)}_{\text{max}}-\big{(}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\big{)}_{\text{max}}\Big{|}\Big{|}$
@@
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{M}\text{Lip}(f_{\text{MLP}_{1}})||\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}||$
$\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{M}||\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}||$
@@
A.6.5 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to EMD
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $X\in\mathcal{S}(\mathbb{R}^{d})$, $\forall X\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. We have shown in subsection A.6.4 above that:
$\bigg{|}\bigg{|}f_{\text{MLP}_{2}}\Big{(}\big{(}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\big{)}_{\text{max}}\Big{)}-f_{\text{MLP}_{2}}\Big{(}\big{(}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\big{)}_{\text{max}}\Big{)}\bigg{|}\bigg{|}$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{M}(X,Y)$
@@ A.6.5 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to EMD
$\bigg{|}\bigg{|}f_{\text{MLP}_{2}}\Big{(}\big{(}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\big{)}_{\text{max}}\Big{)}-f_{\text{MLP}_{2}}\Big{(}\big{(}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\big{)}_{\text{max}}\Big{)}\bigg{|}\bigg{|}$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{M}(X,Y)$
@@
$[\mathbf{F}]_{ij}=\begin{cases}\frac{1}{n(n+1)}&\text{if }j=n+1\\ \frac{1}{n+1}&\text{otherwise}\\ \end{cases}$
@@
$\begin{split}[\mathbf{F}]_{ij}&>0,\qquad 1\leq i\leq n,\quad 1\leq j\leq n+1\\ \sum_{j=1}^{n+1}[\mathbf{F}]_{ij}&=\frac{1}{n},\qquad 1\leq i\leq n\\ \sum_{i=1}^{n}[\mathbf{F}]_{ij}&=\frac{1}{n+1},\qquad 1\leq j\leq n+1\\ \end{split}$
@@
$\sum_{i=1}^{n}\sum_{j=1}^{n+1}[\mathbf{F}]_{ij}\,||\mathbf{v}_{i}-\mathbf{v}_{j}||=\frac{1}{n(n+1)}\sum_{i=1}^{n}||\mathbf{v}_{i}-\mathbf{v}_{n+1}||$
@@
$d_{\text{EMD}}(X,X^{\prime})\leq\frac{1}{n(n+1)}\sum_{i=1}^{n}||\mathbf{v}_{i}-\mathbf{v}_{n+1}||$
@@
For each $i\in[n]$, we have that $\min_{j\in[n+1]}||\mathbf{v}_{i}-\mathbf{v}_{j}||=||\mathbf{v}_{i}-\mathbf{v}_{i}||=0$. Therefore, we have that:
$h(X,X^{\prime})=\max_{i\in[n]}\min_{j\in[n+1]}||\mathbf{v}_{i}-\mathbf{v}_{j}||=0$
@@
$d_{H}(X,X^{\prime})=h(X^{\prime},X)=\max_{i\in[n+1]}\min_{j\in[n]}||\mathbf{v}_{i}-\mathbf{v}_{j}||=0$
For each $i\in[n]$, we have that $\min_{j\in[n]}||\mathbf{v}_{i}-\mathbf{v}_{j}||=||\mathbf{v}_{i}-\mathbf{v}_{i}||=0$. Therefore, we have that:
$d_{H}(X,X^{\prime})=h(X^{\prime},X)=\min_{j\in[n]}||\mathbf{v}_{n+1}-\mathbf{v}_{j}||$
Appendix B Stability of Neural Networks for Sets Under Perturbations
Lemma 3.4 implies that the output variation of $\textsc{NN}_{\textsc{mean}}$, $\textsc{NN}_{\textsc{sum}}$ and $\textsc{NN}_{\textsc{max}}$ under perturbations of the elements of an input set can be bounded via the EMD and Hausdorff distance between the input and perturbed sets, respectively. We next investigate what are the values of EMD, Hausdorff distance and matching distance when an element of a multiset is replaced by a new element.
Proposition B.1.
@@ Appendix B Stability of Neural Networks for Sets Under Perturbations
    $d_{\text{EMD}}(X,X^{\prime})=\frac{1}{n}||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||\qquad\text{and}\qquad d_{\text{M}}(X,X^{\prime})=||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||$
@@
    $d_{\text{H}}(X,X^{\prime})\leq||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||$
@@
$d_{M}(X,X^{\prime})$ $=\min_{\pi\in\mathfrak{S}_{n}}\bigg{[}\sum_{i=1}^{n}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}||\bigg{]}$
$\leq||\mathbf{v}_{1}-\mathbf{v}_{1}||+||\mathbf{v}_{2}-\mathbf{v}_{2}||+\ldots+||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||+\ldots+||\mathbf{v}_{n}-\mathbf{v}_{n}||$
$=||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||$
@@
$||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||$ $=||\mathbf{v}_{1}+\ldots+\mathbf{v}_{i}+\ldots+\mathbf{v}_{n}-\mathbf{v}_{1}-\ldots-\mathbf{v}_{i}^{\prime}-\ldots-\mathbf{v}_{n}||$
$=||\mathbf{v}_{1}+\ldots+\mathbf{v}_{i}+\ldots+\mathbf{v}_{n}-\mathbf{u}_{1}-\ldots-\mathbf{u}_{i}-\ldots-\mathbf{u}_{n}||$
$\leq||\mathbf{v}_{\pi^{*}(1)}-\mathbf{u}_{1}||+\ldots+||\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}||+\ldots+||\mathbf{v}_{\pi^{*}(n)}-\mathbf{u}_{n}||$
@@
We showed that $d_{M}(X,X^{\prime})\leq||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||$ and that $||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||\leq d_{M}(X,X^{\prime})$. Therefore, $d_{M}(X,X^{\prime})=||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||$. Since $|X|=|X^{\prime}|=n$, by Proposition 2.3, we have that $d_{M}(X,X^{\prime})=n\,d_{\text{EMD}}(X,X^{\prime})$. The following then holds:
$d_{\text{EMD}}(X,X^{\prime})=\frac{1}{n}||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||$
@@
For each $j\in[i-1]\cup{i+1,\ldots,n}$, we have that $\min_{k\in[n]}||\mathbf{v}_{j}-\mathbf{u}_{k}||=||\mathbf{v}_{j}-\mathbf{v}_{j}||=0$. For each $j\in[i-1]\cup{i+1,\ldots,n}$, we also have that $\min_{k\in[n]}||\mathbf{u}_{j}-\mathbf{v}_{k}||=||\mathbf{v}_{j}-\mathbf{v}_{j}||=0$. Therefore, we have that:
$\begin{split}h(X,X^{\prime})&=\min_{j\in[n]}||\mathbf{v}_{i}-\mathbf{u}_{j}||\leq||\mathbf{v}_{i}-\mathbf{u}_{i}||=||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||\\ h(X^{\prime},X)&=\min_{j\in[n]}||\mathbf{u}_{i}-\mathbf{v}_{j}||\leq||\mathbf{u}_{i}-\mathbf{v}_{i}||=||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||\\ \end{split}$
Both $h(X,X^{\prime})$ and $h(X^{\prime},X)$ are thus no greater than $||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||$. We then have that:
$\begin{split}d_{H}(X,X^{\prime})&=\max\big{(}h(X,X^{\prime}),h(X^{\prime},X)\big{)}\\ &\leq\max\big{(}||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||,||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||\big{)}\\ &=||\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}||\end{split}$
@@
$||\mathbf{v}_{i}-\mathbf{v}_{i}-\mathbf{u}_{i}||=||\mathbf{u}_{i}||\leq\sqrt{\underbrace{k^{2}+k^{2}+\ldots+k^{2}}_{d\text{ times}}}=k\sqrt{d}$
@@
$d_{M}(X,X^{\prime})$ $=\min_{\pi\in\mathfrak{S}_{n}}\bigg{[}\sum_{i=1}^{n}||\mathbf{v}_{\pi(i)}-\mathbf{v}_{i}-\mathbf{u}_{i}||\bigg{]}$
$\leq||\mathbf{v}_{1}-\mathbf{v}_{1}-\mathbf{u}_{1}||+||\mathbf{v}_{2}-\mathbf{v}_{2}-\mathbf{u}_{2}||+\ldots+||\mathbf{v}_{n}-\mathbf{v}_{n}-\mathbf{u}_{n}||$
$=||-\mathbf{u}_{1}||+||-\mathbf{u}_{2}||+\ldots+||-\mathbf{u}_{n}||$
@@
$\begin{split}h(X,X^{\prime})&=\max_{i\in[n]}\min_{j\in[n]}||\mathbf{v}_{i}-\mathbf{v}_{j}-\mathbf{u}_{j}||\leq\max_{i\in[n]}||\mathbf{v}_{i}-\mathbf{v}_{i}-\mathbf{u}_{i}||=\max_{i\in[n]}||-\mathbf{u}_{i}||\leq k\sqrt{d}\\ h(X^{\prime},X)&=\max_{i\in[n]}\min_{j\in[n]}||\mathbf{v}_{i}+\mathbf{u}_{i}-\mathbf{v}_{j}||\leq\max_{i\in[n]}||\mathbf{v}_{i}+\mathbf{u}_{i}-\mathbf{v}_{i}||=\max_{i\in[n]}||\mathbf{u}_{i}||\leq k\sqrt{d}\\ \end{split}$
@@
Appendix C Experimental Setup

### v3

@@ B.8.3 The $\textsc{NN}_{\textsc{sum}}$ model is Lipschitz continuous with respect to EMD
$\bigg\|f_{\text{MLP}_{2}}\bigg(\sum_{i=1}^{M}f_{\text{MLP}_{1}}(\mathbf{v}_{i})\bigg)-f_{\text{MLP}_{2}}\bigg(\sum_{j=1}^{M}f_{\text{MLP}_{1}}(\mathbf{u}_{j})\bigg)\bigg\|$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{M}(X,Y)$
@@
B.8.4 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to the matching distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. Let $f_{\text{MLP}_{1}}(X)=\big\{\!\!\big\{f_{\text{MLP}_{1}}(\mathbf{v}_{1}),\ldots,f_{\text{MLP}_{1}}(\mathbf{v}_{m})\big\}\!\!\big\}$ and $f_{\text{MLP}_{1}}(Y)=\big\{\!\!\big\{f_{\text{MLP}_{1}}(\mathbf{u}_{1}),\ldots,f_{\text{MLP}_{1}}(\mathbf{u}_{n})\big\}\!\!\big\}$. Let also $\big(f_{\text{MLP}_{1}}(X)\big)_{\text{max}}$ and $\big(f_{\text{MLP}_{1}}(Y)\big)_{\text{max}}$ denote the output of the max operator applied to multisets $f_{\text{MLP}_{1}}(X)$ and $f_{\text{MLP}_{1}}(Y)$, respectively. Let $\pi^{*}$ denote the matching produced by the solution of the matching distance function:
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{M}}\sum_{i=1}^{M}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}\|\\$
@@ B.8.4 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to the matching distance
$\bigg\|f_{\text{MLP}_{2}}\Big(\big(f_{\text{MLP}_{1}}(X)\big)_{\text{max}}\Big)-f_{\text{MLP}_{2}}\Big(\big(f_{\text{MLP}_{1}}(Y)\big)_{\text{max}}\Big)\bigg\|$ $\leq\text{Lip}(f_{\text{MLP}_{2}})\Big\|\big(f_{\text{MLP}_{1}}(X)\big)_{\text{max}}-\big(f_{\text{MLP}_{1}}(Y)\big)_{\text{max}}\Big\|$
@@
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{M}\Big\|f_{\text{MLP}_{1}}(\mathbf{v}_{\pi^{*}(i)})-f_{\text{MLP}_{1}}(\mathbf{u}_{i})\Big\|$
$\leq\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{M}\text{Lip}(f_{\text{MLP}_{1}})\|\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}\|$
$\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})\sum_{i=1}^{M}\|\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}\|$
@@
B.8.5 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to EMD
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. We have shown in subsection B.8.4 above that:
$\bigg\|f_{\text{MLP}_{2}}\Big(\big(f_{\text{MLP}_{1}}(\mathbf{v}_{i})\big)_{\text{max}}\Big)-f_{\text{MLP}_{2}}\Big(\big(f_{\text{MLP}_{1}}(\mathbf{u}_{j})\big)_{\text{max}}\Big)\bigg\|$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{M}(X,Y)$
@@ B.8.5 The $\textsc{NN}_{\textsc{max}}$ model is Lipschitz continuous with respect to EMD
$\bigg\|f_{\text{MLP}_{2}}\Big(\big(f_{\text{MLP}_{1}}(\mathbf{v}_{i})\big)_{\text{max}}\Big)-f_{\text{MLP}_{2}}\Big(\big(f_{\text{MLP}_{1}}(\mathbf{u}_{j})\big)_{\text{max}}\Big)\bigg\|$ $\leq\text{Lip}(f_{\text{MLP}_{1}})\text{Lip}(f_{\text{MLP}_{2}})d_{M}(X,Y)$
@@
[…]
@@
[…]
@@
$\sum_{i=1}^{n}\sum_{j=1}^{n+1}[\mathbf{F}]_{ij}\,\|\mathbf{v}_{i}-\mathbf{v}_{j}\|=\frac{1}{n(n+1)}\sum_{i=1}^{n}\|\mathbf{v}_{i}-\mathbf{v}_{n+1}\|$
@@
$d_{\text{EMD}}(X,X^{\prime})\leq\frac{1}{n(n+1)}\sum_{i=1}^{n}\|\mathbf{v}_{i}-\mathbf{v}_{n+1}\|$
@@
For each $i\in[n]$, we have that $\min_{j\in[n+1]}\|\mathbf{v}_{i}-\mathbf{v}_{j}\|=\|\mathbf{v}_{i}-\mathbf{v}_{i}\|=0$. Therefore, we have that:
$h(X,X^{\prime})=\max_{i\in[n]}\min_{j\in[n+1]}\|\mathbf{v}_{i}-\mathbf{v}_{j}\|=0$
@@
$d_{H}(X,X^{\prime})=h(X^{\prime},X)=\max_{i\in[n+1]}\min_{j\in[n]}\|\mathbf{v}_{i}-\mathbf{v}_{j}\|$
For each $i\in[n]$, we have that $\min_{j\in[n]}\|\mathbf{v}_{i}-\mathbf{v}_{j}\|=\|\mathbf{v}_{i}-\mathbf{v}_{i}\|=0$. Therefore, we have that:
$d_{H}(X,X^{\prime})=h(X^{\prime},X)=\min_{j\in[n]}\|\mathbf{v}_{n+1}-\mathbf{v}_{j}\|$
Appendix C Stability of Neural Networks for Sets Under Perturbations
Lemma 3.5 implies that the output variation of $\textsc{NN}_{\textsc{mean}}$, $\textsc{NN}_{\textsc{sum}}$ and $\textsc{NN}_{\textsc{max}}$ under perturbations of the elements of an input set can be bounded via the EMD, matching distance and Hausdorff distance between the input and perturbed sets, respectively. We next investigate what are the values of EMD, matching distance and Hausdorff distance when an element of a multiset is replaced by a new element.
Proposition C.1.
@@ Appendix C Stability of Neural Networks for Sets Under Perturbations
    $d_{\text{EMD}}(X,X^{\prime})=\frac{1}{n}\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|\qquad\text{and}\qquad d_{\text{M}}(X,X^{\prime})=\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|$
@@
    $d_{\text{H}}(X,X^{\prime})\leq\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|$
@@
$d_{M}(X,X^{\prime})$ $=\min_{\pi\in\mathfrak{S}_{n}}\bigg[\sum_{i=1}^{n}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}\|\bigg]$
$\leq\|\mathbf{v}_{1}-\mathbf{v}_{1}\|+\|\mathbf{v}_{2}-\mathbf{v}_{2}\|+\ldots+\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|+\ldots+\|\mathbf{v}_{n}-\mathbf{v}_{n}\|$
$=\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|$
@@
$\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|$ $=\|\mathbf{v}_{1}+\ldots+\mathbf{v}_{i}+\ldots+\mathbf{v}_{n}-\mathbf{v}_{1}-\ldots-\mathbf{v}_{i}^{\prime}-\ldots-\mathbf{v}_{n}\|$
$=\|\mathbf{v}_{1}+\ldots+\mathbf{v}_{i}+\ldots+\mathbf{v}_{n}-\mathbf{u}_{1}-\ldots-\mathbf{u}_{i}-\ldots-\mathbf{u}_{n}\|$
$\leq\|\mathbf{v}_{\pi^{*}(1)}-\mathbf{u}_{1}\|+\ldots+\|\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}\|+\ldots+\|\mathbf{v}_{\pi^{*}(n)}-\mathbf{u}_{n}\|$
@@
We showed that $d_{M}(X,X^{\prime})\leq\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|$ and that $\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|\leq d_{M}(X,X^{\prime})$. Therefore, $d_{M}(X,X^{\prime})=\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|$. Since $|X|=|X^{\prime}|=n$, by Proposition 2.3, we have that $d_{M}(X,X^{\prime})=n\,d_{\text{EMD}}(X,X^{\prime})$. The following then holds:
$d_{\text{EMD}}(X,X^{\prime})=\frac{1}{n}\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|$
@@
For each $j\in[i-1]\cup\{i+1,\ldots,n\}$, we have that $\min_{k\in[n]}\|\mathbf{v}_{j}-\mathbf{u}_{k}\|=\|\mathbf{v}_{j}-\mathbf{v}_{j}\|=0$. For each $j\in[i-1]\cup\{i+1,\ldots,n\}$, we also have that $\min_{k\in[n]}\|\mathbf{u}_{j}-\mathbf{v}_{k}\|=\|\mathbf{v}_{j}-\mathbf{v}_{j}\|=0$. Therefore, we have that:
[…]
Both $h(X,X^{\prime})$ and $h(X^{\prime},X)$ are thus no greater than $\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|$. We then have that:
$\begin{split}d_{H}(X,X^{\prime})&=\max\big(h(X,X^{\prime}),h(X^{\prime},X)\big)\\ &\leq\max\big(\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|,\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|\big)\\ &=\|\mathbf{v}_{i}-\mathbf{v}_{i}^{\prime}\|\end{split}$
@@
$\|\mathbf{v}_{i}-\mathbf{v}_{i}-\mathbf{u}_{i}\|=\|\mathbf{u}_{i}\|\leq\sqrt{\underbrace{k^{2}+k^{2}+\ldots+k^{2}}_{d\text{ times}}}=k\sqrt{d}$
@@
$d_{M}(X,X^{\prime})$ $=\min_{\pi\in\mathfrak{S}_{n}}\bigg[\sum_{i=1}^{n}\|\mathbf{v}_{\pi(i)}-\mathbf{v}_{i}-\mathbf{u}_{i}\|\bigg]$
$\leq\|\mathbf{v}_{1}-\mathbf{v}_{1}-\mathbf{u}_{1}\|+\|\mathbf{v}_{2}-\mathbf{v}_{2}-\mathbf{u}_{2}\|+\ldots+\|\mathbf{v}_{n}-\mathbf{v}_{n}-\mathbf{u}_{n}\|$
$=\|-\mathbf{u}_{1}\|+\|-\mathbf{u}_{2}\|+\ldots+\|-\mathbf{u}_{n}\|$
@@
[…]
@@
Appendix D Experimental Setup